OpenStudy (jessicah123):

Write the rate and unit rate that represents the following situations. 3 baskets in 45 minutes 

OpenStudy (jessicah123):

@Directrix

OpenStudy (steve816):

The rate would be\[\huge \frac{ 3~baskets }{ 45~minutes }\] Which can be simplified to the unit rate of \[\huge \frac{ 1~basket }{ 15~minutes }\]

OpenStudy (steve816):

Welcome to OpenStudy!!!!!!!!!!!!!!!!!!!!!

OpenStudy (jessicah123):

but isn't unit rate supposed to have a 1 as a denomintor?

OpenStudy (jessicah123):

@steve816

OpenStudy (mathstudent55):

\(\dfrac{3 ~baskets}{45~minutes}= \dfrac{1 ~basket}{15~minutes} = \dfrac{1}{15}~\dfrac{baskets}{minute} \)

OpenStudy (jessicah123):

but isn't unit rate supposed to have a 1 as a denomintor?

OpenStudy (mathstudent55):

baskets/minute is a unit rate because it is per 1 minute.

OpenStudy (mathstudent55):

1 basket per 15 minutes is not a unit rate, but 1/15 baskets per minute is.

OpenStudy (jessicah123):

I am very confused